### 07 Dec 2023 v1

# A general framework for the generation of probabilistic socioeconomic scenarios: Quantification of national food security risk with application to the cases of Egypt and Ethiopia

Abstract
In this work a general framework for providing detailed probabilistic socioeconomic scenarios as well as estimates concerning country-level food security risk is proposed. Our methodology builds on (a) the Bayesian probabilistic version of the world population model and (b) on the interdependencies of the minimum food requirements and the national food system capacities on key drivers, such as: population, income, natural resources, and other socioeconomic and climate indicators. Model uncertainty plays an important role in such endeavours. In this perspective, the concept of the recently developed convex risk measures which mitigate the model uncertainty effects, is employed for the development of a framework for assessment, in the context of food security. The proposed method provides predictions and evaluations for food security risk both within and across probabilistic scenarios at country level. Our methodology is illustrated through its implementation for the cases of Egypt and Ethiopia, for the time period 2019-2050, under the combined context of the Shared Socioeconomic Pathways (SSPs) and the Representative Concentration Pathways (RCPs).
Keywords: food security risk; model uncertainty; probabilistic projections; risk quantification; Representative Concentration Pathways; Shared Socioeconomic Pathways;
1 Introduction & Basic Concepts on Food Security
National food security is a key issue in sustainability studies, while a key factor in such studies is population Garnett (2013); Premanandh (2011). According to World Food Summit Summit (1996):
”Food security exists when all people, at all times, have physical and economic access to sufficient, safe and nutritious food that meets their dietary needs and food preferences for an active and healthy life”.
Under this definition, four different dimensions [^2] https://www.worldbank.org/en/topic/agriculture/brief/food-security-update/what-is-food-security are covered: (a) physical availability of food, (b) economic and physical access to food, (c) food utilization and (d) stability (sustainability) of the other three dimensions over time. In this work, we focus on the modelling, quantification and prediction of food security risk at country level, following a bottom-up probabilistic modelling approach relying mainly on a key factor that affects food security, i.e. population and its structure. Probabilistic population projections are generated under plausible socioeconomic scenarios and in particular under the so called Shared Socioeconomic Pathways (SSP) framework Lutz et al. (2017); Lutz et al. (2018), and then are incorporated to a global macroeconomic model for generating SSP-compatible projections for key drivers of national economies, like GDP, labour, etc. Then, these probabilistic modelling approaches, accompanied by standard econometric models for describing certain components that affect food security and taking into account the effect of climate drivers exogenously, are used to construct a decision support tool for assessing national food security risk. We insist on the term national food security since at a different scale, e.g. at household level, the situation could differ significantly [^3] Australia, Canada, UK and USA display high food security at national level but significant food insecurity at household level (see e.g. Loopstra (2018))). Therefore, in the perspective that food security risk is treated in this work, two main quantities of interest are considered as the key components: (a) the minimum food requirements of the population at a specific point in time (in terms of caloric content, rather than recommended dietary composition), and (b) the national food system capacity at the same time, i.e. the total quantity of the food (again in terms of caloric content) that is available to the country’s population. Note that finer aspects like dietary needs, food habits and poverty issues that may restrict access to food, are not directly treated in our approach. However, the proposed framework, can be extended to account for alternative modelling considerations, hence issues such as the above can be introduced if required. In what follows, by the term food requirements we will mean caloric requirements or intake.
Minimum food requirements for subsistence [^4] Note that the minimum caloric intake for bare survival does not necessarily coincide with the minimum calorie requirements that guarantee nutritional security depend in an inelastic way on population, as well as on its detailed age structure and each age group’s activity levels. On the other hand, food self-sufficiency at national level is mainly determined by the capacity of the country’s food system, indicated by the food amount (calories) available for human consumption as a result of the system functioning. This term is very accurately described in the definition provided by Food and Agriculture Organization (FAO) [^5] FAO (1999) https://www.fao.org/3/X3936E/X3936E00.htm:
“The concept of food self-sufficiency is generally taken to mean the extent to which a country can satisfy its food needs from its own domestic production”.
The capacity of the food system is importantly affected by the country’s economy, through the production system that involves planning of the economic sectors involved in food production and/or distribution, choice of crops, imports, exports, etc.
Reliable predictions for both minimum caloric requirements and food system capacity, in a long-time horizon, are very important in determining future food balance. Quantifying food security risk on a sufficiently long horizon, provides policy makers the luxury of adopting long-term measures, combining a portfolio of production restructuring policies, international trade treaties, adoption of scientific measures or modern technologies etc, which may efficiently alleviate the risk of future food security. Such trustworthy predictions will inevitably rely upon driver trends, and must be probabilistic in nature, i.e. providing the distribution of the relevant random factors, giving full information of the trends along with their validity, rather than point estimates which carry less information concerning the predictions. They should also take into account model uncertainty, which is inherent in these approaches, especially when long-term predictions are involved, because the effects of stochastic factors on these predictions may be at best partially known.
In the relevant literature, much attention has been given to the impacts of climate change in food security risk evaluations (see e.g. Schmidhuber and Tubiello (2007) and references therein). Discussions concerning the importance for the determination of measures and the actions to be taken in order to counter food security challenges due to climate change and extreme climate events are provided in Campbell et al. (2016) and Hasegawa et al. (2021). However, Hasegawa et al. (2018) argue that stringent climate mitigation policies would have a greater negative impact on global hunger and food consumption than the direct impacts of climate change. From the modelling perspective of food security, several directions are discussed in the literature introducing methodological frameworks (see e.g. Krishnamurthy et al. (2014)) and recognizing the need to better model key factors of food demand like demographics and economic growth due to the uncertainty that are subject to Valin et al. (2014). It seems that scenario-based approaches Hasegawa et al. (2015); Molotoks et al. (2021); Van Dijk et al. (2021) and attempts employing global models van Meijl et al. (2020) or integrating different models Müller et al. (2020), lead to more accurate assessments of food security risk, while simultaneously treating the model uncertainty issues. Although the majority of papers study the evaluation of the global food security risk, there are some works that study the matter on regional or country level Chen et al. (2021); Mainuddin and Kirby (2015) following the aforementioned approaches. In this work, a scenario-based approach is adopted for evaluating food security risk at country level, considering various socio-economic and climate scenarios under the probabilistic setting (i.e. each scenario is considered as a probability model which can produce a number of different trajectories) which to the best of our knowledge has not been yet proposed in the relevant literature. To this direction, the typical framework of the Shared Socioeconomic Pathways is revised to be compatible with this probabilistic setting, and then relevant modelling approaches are followed. Moreover, a food security risk index which is able to discount food security risk within and across the considered probabilistic socio-economic combined with climate scenarios is proposed.
More specifically, we provide a suitable framework for assessing food security risk at country level, following a bottom-up modelling approach, building on the probabilistic extension of the population model introduced in Raftery et al. (2012). The underlying population model relies on a Bayesian hierarchical modelling (BHM) procedure (see e.g. Gelman et al. (1995)) and provides detailed information on the future population of the world and its age structure by country under the probabilistic scenarios framework, i.e. for a certain time instant a distribution of the possible population counts is provided instead of a single point prediction. Our first contribution, is that by an appropriate modification of the structure of the population model, we are allowed to generate probabilistic population scenarios under the SSP framework, i.e. generating population trajectories of a priori known SSP status. Our second contribution is the incorporation of the generated population projections into a global macroeconomic model (in our case MaGE Fouré et al. (2013); Fontagné et al. (2022)) to generate SSP-compatible future scenarios for important drivers of economy like GDP, labour, etc. Finally, this probabilistic scenarios generation scheme is combined with climate scenarios, within the context of the Representative Concentrations Pathways (RCP) (see e.g. Meinshausen et al. (2020)), thus providing a complete framework for policies assessment for taking into account uncertainty on future predictions. Importantly, the probabilistic modelling framework adopted in this work, allows us to turn the qualitative narratives of the SSP/RCP scenarios into concrete quantitative scenarios, thus turning them into important quantitative policy and decision making tools (see Section 2.2).
The proposed framework is then implemented in quantifying food security risk at country level, for both between and across a selected set of combined socioeconomic and climate scenarios (SSP-RCP), under the aforementioned probabilistic setting. In this direction, the following modelling considerations are performed. First, employing the projected population structures (by age and gender), as well as the minimum required calories intake per age group and gender (determined by nutrition experts), estimations are provided for the minimum calories intake required for subsistence for the population at national level. Next, the national food system capacities under their current food system patterns are estimated, building models relying on basic socioeconomic drivers (population, income, labour) and climate factors (temperature, precipitation). These structural models of food system capacities are calibrated on past data and then, incorporating the population projections, as well as those of other related socioeconomic drivers, probabilistic projections on food system capacities are provided for a time period of interest (in our case 2019-2050). Combining the above modelling approaches, a new food security risk indicator is proposed. The advantages of the new index is that (a) it requires less detailed data for its calculation compared to well established food security risk indicators (please see the relevant discussion at the beginning of Section 3), (b) it is more affordable in terms of computational cost, (c) it is able to project the food security situation in the future, and (c) it takes into account natural resources, such as for example the pressure on national water sources. The construction of this index relies on the concept of the convex risk measures (see e.g. Detlefsen and Scandolo (2005); Föllmer and Schied (2002); Frittelli and Gianin (2002)) allowing for mitigating successfully the effects of model uncertainty and providing robust risk estimates either within a certain probabilistic scenario or across different scenarios. The latter risk estimates rely on the concept of Fréchet utilities or risk measures (Papayiannis and Yannacopoulos (2018), Petracou et al. (2022)) and are crucial, especially for long-term effects and policies, that have to be predicted and implemented long before the actual scenario that materializes has been fully clarified.
The paper is organized in the following manner: In Section 2 the probabilistic extension of the population model and its modification so that it becomes compatible with the SSP framework and the set of socio-economic and climate scenarios that are considered throughout the paper are introduced. Next, in Section 3 the modelling approaches and considerations for the minimum caloric intake and the food system capacity at country level are presented. Moreover, the new food security risk index is introduced, accompanied by a discussion and comparisons to relevant well established indicators in the literature. Finally, in Section LABEL:sec-3, the proposed approach is implemented for the cases of Egypt and Ethiopia using as training data records from the period 1990-2018 and providing projections till the year 2050 both for within and across the SSP-RCP scenarios framework.
2 A general probabilistic socioeconomic modelling framework for food security
2.1 The main probabilistic population model
Population growth and evolution of the population’s age structure are key factors driving many socioeconomic indices, including economic growth, production, environmental issues, food and water demand etc. In this respect, demographics must be the starting point for any socioeconomic modelling study. In this section, some key results concerning scenario development for future population growth are presented. The state of the art model concerning world population is the probabilistic model proposed by Raftery and coworkers Raftery et al. (2012). This model takes into account the inherent uncertainty of the phenomenon and its effects on future population projections using a Bayesian hierarchical model (see e.g. Congdon (2010); Gelman et al. (1995)). The model is based on the natural evolution of the population phenomenon as characterized by the standard (deterministic) model employed by the United Nations (UN),
(1) $P_{c}(t)-P_{c}(t-1)=B_{c}(t)-D_{c}(t)+M_{c}(t)$
where $P_{c}(t)$ denotes the population of country $c$ at time $t$ (corresponding either to a single year or a 5-year period), $B_{c}(t)$ stands for the number of births (which depends on the total fertility rate of the country), $D_{c}(t)$ denotes the number of deaths (which depends on the life expectancy) and $M_{c}(t)$ measures the net international migration. Uncertainty is introduced to the population model since its main components (fecundity, mortality, migration) are subject to random factors that cannot be sufficiently modelled. The Bayesian approach proposed in Raftery et al. (2012) captures uncertainty on each one of the major components of population through the construction/introduction of distinct hierarchical models for important components such as fertility, mortality and migration [^6] Note that only international migration is considered and not internal migration phenomena like urbanization, and then propagating uncertainty to the output of model (1), i.e. providing probabilistic estimates either for $P_{c}(t)$ or its breakdown into age groups and sex at various future times $t$. Clearly, uncertainty becomes higher as the time progresses. Based on an extensive database of past world population data (recorded population pyramids, fertility rates, mortality rates, etc), the fundamental law (1), and the principles of Bayesian statistics, the probabilistic features of the uncertainty factors driving the population fluctuations are recovered. Then, using this information, the fundamental law (1) is iterated forward and used to obtain estimates for the future evolution of the quantities of interest. The estimates incorporate in a dynamically consistent fashion the effects of ambiguity as documented at least from the past data, and thus provide uncertainty consistent predictions for the future.
One of the key features of the model is that it allows for quantities related to population projections to be random variables, characterized by a probability distribution, rather than point estimates. In particular, instead of producing a point estimate for a population related quantity $X(t)$ at time $t$ ($X$ can represent for example population for a particular age group or sex, or quantities such as fertility, life expectancy, etc.), the model treats $X(t)$ as a random variable and produces (dynamically) a set of possible realizations $\{X^{(j)}(t)\,\,:\,\,j=1,\ldots,n\}$, which are approximations for the probability distribution of $X(t)$, based on possible outcomes of the uncertainty factors driving the phenomenon. Using this probability distribution, one characterizes the quantity $X(t)$ with quantities carrying more information than a mere point estimate, for example its percentiles at certain confidence levels or conditional means. These different realizations
$\Pi:=\{X^{j}(t)\,\,:\,\,t=T_{0},\ldots,T,\,\,j=1,\ldots,n\},$
where $T_{0}<T$ are two selected time horizons, will be referred to as trajectories with $\{X^{j}(t)\,\,:\,\,t=T_{0},\ldots,T\}$ for fixed $j$ representing a particular realization (i.e. a particular possible path) for the evolution of population parameter in the future time interval $[T_{0},T]$. Clearly, only one of the above paths in $\Pi$, if any, will materialize. However, the set of paths $\Pi$ provides us with important information concerning the probability of occurrence of paths with certain characteristics and allows for prediction of future population trends as well as the formulation of scenarios concerning these trends.
Some information on the structure of the probabilistic population model must be introduced here in order to make the SSP scenario generation procedure described in Section 2.2 more clear. In particular, Raftery et al. (2012) rely on model (1), however treating the $B_{c}(t)$ and $D_{c}(t)$ components separately, according to the probabilistic modeling approach mentioned above. First, a hierarchical model is constructed for the Total Fertility Rate (TFR) component which provides projections for the fertility rates distribution at the country level and then for the number of births distribution according to the approach presented in Alkema et al. (2011). Then, this information is used to an hierarchical model for the Life Expectancy (e0) component according to the approach presented in Raftery et al. (2013), which is then used to provide projections for life expectancy distributions of females and males per country at the various age-groups as well as to provide the mortality rates distribution on each age-group by gender. In particular, the life expectancy ($e_{0}$) and the total fertility rate (TFR) components of the model are captured by the parametric equations
(4) $\begin{array}[]{ll}e^{f}_{0,c,t+1}=e^{f}_{0,c,t}+g_{1}(e^{f}_{0,c,t}\mid{\bm{\theta}}^{c}_{1})+\eta^{e_{0}}_{c,t+1},&\eta^{e_{0}}_{c,t+1}\sim N\left(0,\varphi_{1}(e_{0,c,t}^{f})\right)\\ TFR_{c,t+1}=TFR_{c,t+1}+g_{2}(TFR_{c,t}\mid{\bm{\theta}}^{c}_{2})+\eta^{TFR}_{c,t+1},&\eta^{TFR}_{c,t+1}\sim N\left(0,\varphi_{2}(TFR_{c,t})\right)\\ \end{array}$
where the functions $g_{1}(\cdot),g_{2}(\cdot)$ determine separate double logistic-type growth models with respect to $e_{0}^{f}$ and $TFR$ expanding the UN modelling approach to a probabilistic setting (please see Alkema et al. (2011); Raftery et al. (2013) for details), respectively, while the functions $\varphi_{1}(\cdot),\varphi_{2}(\cdot)$ determine the variance terms concerning the residuals of each model. Note, that in system (4) the life expectancy model concerns only the females, while the life expectancy for the males ($e_{0}^{m}$) is obtained by building a gap model with respect to the $e_{0}^{f}$ term, according to the approach described in Raftery et al. (2013). Moreover, the whole modelling approach estimates a set of country-specific parameters (${\bm{\theta}}^{(c)}:=({\bm{\theta}}_{1}^{(c)},{\bm{\theta}}_{2}^{(c)})^{\prime}$) referring to the special characteristics of each country (as displayed by the available data) concerning life expectancy and total fertility rate, while each set ${\bm{\theta}}^{(c)}$ comes as a sample of a world distribution subject to some world-specific parameters (as obtained from the whole training dataset for the population model). In this perspective, the noise introduced in the population projections is carefully parameterized by the empirical evidence both on the total available dataset (world data) and on the characteristics displayed on a local level (country-specific attributes).
Figure 1: The procedure followed for generating probabilistic population projections illustrating the data providers, the separate model components and the respective R packages that are used.
Concerning the third component that contributes to the population model, the net migration (MIG) term (at country level), projections for the future states are collected by the UN and other databases (see e.g. Wittgenstein Centre Database [^7] http://www.wittgensteincentre.org) and then incorporated to the main population model (1). Note that although there are similar hierarchical modeling approaches for migration in the literature (e.g. Azose and Raftery (2015)) as the ones discussed above for the other two components, the lack or insufficiency of migration data for all the countries in our area of interest makes the implementation of this model infeasible at present, so we resort to the simple modelling framework mentioned in the previous sentence. Finally, all the above components are combined and aggregated in the general population model (1) to provide the future population projections in terms of trajectories (possible scenarios) or distributions if conditioned to certain time instants. All the above modeling task is implemented in the statistical software R through the related package bayesPop described in detail in Ševčíková and Raftery (2016). The roadmap of the whole modeling task is illustrated in Figure (1).
2.2 Shared Socioeconomic Pathways, Population and GDP Projections
The concept of scenario making concerning future events has infiltrated environmental economics and has become a fundamental tool in the analysis of potential future outcomes. An important class of scenarios used frequently in analyses are the Shared Socioeconomic Pathways O’Neill et al. (2014), which set certain plausible assumptions concerning the evolution for key socioeconomic drivers (such as population growth or fertility) for certain parts of the world. The possible states of the world are divided into five qualitative scenarios (rapid, medium, stalled, inequality and development) according to the levels of specific demographic characteristics and specifically, fertility, life expectancy (or mortality), migration and education. Since the adopted population projection method does not take into account the education levels, this factor is omitted for the purposes of this work.
These narratives are phrased in a qualitative fashion (please see Table 1), so they have to be transcribed to quantitative terms if they are to be used constructively in quantitative models. For instance, under the SSP1 scenario presented in Table 1 the evolution of life expectancy in high fertility countries is characterised as high, with the obvious question arising being: “How high is high? Which figure for the life expectancy in a particular country can safely fit (i.e. with a particular probability) to the qualitative characterization high? An important contribution of our work, and a key point in our approach is the transcription of the SSPs in a quantitative, probabilistic setting. This allows us to provide the population projections (appropriately quantified in terms of probability distributions or sample paths for the quantities of interest) under each one of these pathways.
Socio-Economic Scenario Country Fertility Life Migration
grouping expectancy
HiFert Low High Medium
SSP1: Sustainability LoFert Low High Medium
Rich-OECD Medium High Medium
HiFert Medium Medium Medium
SSP2: Middle of the road LoFert Medium Medium Medium
Rich-OECD Medium Medium Medium
HiFert High Low Low
SSP3: Fragmentation LoFert High Low Low
Rich-OECD Low Low Low
HiFert High Low Medium
SSP4: Inequality LoFert Low Medium Medium
Rich-OECD Low Medium Medium
HiFert Low High High
SSP5: Conventional development LoFert Low High High
Rich-OECD High High High
Table 1: The Shared Socioeconomic Pathways (SSP) definitions
SSP scenarios may differ for each country, depending on its grouping as (a) high fertility country (HiFert), (b) low fertility country (LoFert) or (c) Rich-OECD country [^8] http://www.oecd.org/about/members-and-partners/. Specifications of the various SSPs with respect to the country groupings, as described in Lutz et al. (2017); Lutz et al. (2018), are illustrated in Table 1. As already mentioned, serious drawback of defining the scenarios in this fashion, is the difficulty to adopt a common line on what is meant by low, medium and high for the various quantities of interest. This is of paramount importance in quantitative modelling, upon which policy making will be based. For instance, assume that we need to quantify the low fertility sub-scenario for both a LoFert and a HiFert country. According to the UN fertility modelling the value 2.1 is considered as the threshold value under which the population of a country declines and above which the population increases.
If we are planning to generate population scenarios for a relative short time period (e.g. 20-30 years from now), according to the aforementioned threshold for TFR, we might not be able to generate all possible TFR sub-scenarios for all countries, due to the different dynamics of the collected data. In this perspective, for a LoFert country an SSP scenario that includes ”high TFR” may not be feasible, while not feasible might be also for a HiFert country an SPP scenario that includes ”low TFR”. Therefore, this universal determination for high-low cases leads to the problematic situation that a policy maker might not be able to create all the SSP scenarios for all countries at any time horizon, a fact that can cause problems in environmental modelling. However, as shown in this section, the probabilistic approach to modeling (such as for example the population model described in Section 2.1) can be nicely integrated with SSP qualitative scenarios turning them into well defined quantitative tools. This, provides a concrete and realistic framework for scenario building, in which the various sub-scenarios (Low, Medium and High) are endogenously and consistently determined by the evolutionary dynamics of the system under study.
The quantification of the SSP scenarios can be effected as follows: Using the probabilistic population model presented in Section 2.1, samples of trajectories for the possible realizations of the evolution over time for the fertility and life expectancy components for each country of interest are generated. These samples are used to obtain the evolution of the probability distribution of these quantities over time, and the relevant quantiles of these distributions at specific time instances are then used to determine the quantitative thresholds that characterise the sub-scenarios Low, Medium and High, as mentioned in Table 1. For instance, if we are interested in determining the range of the three different sub-scenarios (Low, Medium and High) for the TFR of a certain country $c$ up to a certain time horizon $T$ (e.g. $T=2050$), we operate as follows: Employing the distribution of the generated trajectories at the time instant $T$ for the particular country of interest $c$, we separate the induced distribution in three equal parts using the 33% and 66% quantiles, and then use these values as lower and upper thresholds to distinguish the sample of trajectories to the three different sub-scenarios. According to this rule, a trajectory is assigned to the high sub-scenario if at the terminal year 2050 the observation for the corresponding quantity lies on the top 66% of the empirical distribution (i.e. if for trajectory $j$ holds that $TFR_{c,2050,j}\geq q_{TFR,c,2050}(0.66)$ with $q_{TFR,c,2050}(\cdot)$ denoting the induced from the generated sample quantile function for TFR at year $2050$ for the country $c$). The classification of trajectories to the other sub-scenarios is performed in a similar manner.
Population Driver Sub-Scenario Threshold Value
Lower Upper
Low $q_{TFR,c,t=T}(0.33)$
Total Fertility Rate (TFR) Medium $q_{TFR,c,t=T}(0.33)$ $q_{TFR,c,t=T}(0.66)$
High $q_{TFR,c,t=T}(0.66)$
Low $q_{e_{0},c,t=T}(0.33)$
Life Expectancy ($e_{0}$) Medium $q_{e_{0},c,t=T}(0.33)$ $q_{e_{0},c,t=T}(0.66)$
High $q_{e_{0},c,t=T}(0.66)$
Low As specified in Lutz et al. (2018) (deterministic)
Net Migration (MIG) Medium As specified in Lutz et al. (2018) (deterministic)
High As specified in Lutz et al. (2018) (deterministic)
Table 2: Population drivers sub-scenarios definitions based on the generated sample of trajectories
One advantage of this methodology is that the corresponding levels for the scenarios are not preassigned, as for example in the UN methodology, but are determined endogenously by the history and dynamics of the data from the population model 999which by its Bayesian nature is calibrated using a vast global database of past population data, hence carries important information on the history of the population phenomenon.. Repeating the procedure above for each of the trajectories in the sample for each country, we end up with three sub-samples (one triplet for fertility rate and one triplet for life expectancy resp.) each corresponding to different possible realizations of the low, medium and high intensity levels (sub-scenarios). These sub-samples provide important statistical information, such as moments, variability, etc. within scenarios (i.e. within the Low, Medium and High sub-scenarios described in Table 1), and can be used to generate further scenarios and the calculation of conditional expected values of other quantities of interest, depending on the fundamental quantities (i.e. TFR or $e_{0}$ etc) modeled by these trajectories. A possible question to this procedure could be concerning the validity of the generated sample of trajectories by the Bayesian model for the key population drivers (TFR, $e_{0}$) and its ability to represent all possible (realistic) future states of the world. Since the Bayesian model relies on the observed data from previous periods and takes into account possible relations of the country or region under study with all other countries and regions of the world, then any reasonable scenario (with respect to the data that have been collected up to the time the projection task is executed) should be amenable to simulation. Using a sufficiently large number of simulated trajectories, e.g. 100,000 trajectories, should guarantee that the results are reliable. The country-specific discrimination rule for each population driver is illustrated in Table 2. Note that migration levels could be also determined in the same manner if a probabilistic approach had been used (see e.g. Azose and Raftery (2015); Azose et al. (2016)), however in our case for simplicity, and since for the application we consider migration will not play a major role we use the pointwise net migration projections under each SSP scenario, as provided by the Wittgenstein Center database.
Figure 2: Median probabilistic population projections under each SSP scenario for Egypt and Ethiopia.
We follow this approach for each country of interest and generate a sufficient database of possible future realizations of the population phenomenon, as obtained by model (1), which would be compatible with the various SSP scenarios, and the associated sub-scenarios on the population key quantities. For example, for a country belonging to the HiFert group (according to Table 1), the SSP2 scenario consists of the TFR trajectories that belong to the Medium TFR sub-scenario, the e0 trajectories that belong to the Medium e0 sub-scenario and the net migration projection under the medium sub-scenario. These three components are integrated by the population model (1) and provide the population trajectories out of which the SSP2 scenario for this country consists of. This is done in a probabilistic manner, in the sense that we are able to compute quantiles, moments, etc. for the certain population dimensions for this country. In Figure 2 the resulting median population projections for Egypt (EGY) and Ethiopia (ETH) are illustrated under each SSP scenario. Moreover, the population projections in terms of relative growth rates including also the 90% uncertainty zone for both countries in each scenario are illustrated in Figure 10 in Appendix A, while the instantaneous projected total population distributions at year 2050 for Egypt and Ethiopia under all SSP scenarios are illustrated in Appendix B.
Building scenarios representing different pathways for the population and its age structure evolution, offers a vehicle for the estimation of key economic drivers, like labour force, gross domestic product (GDP) and others (see e.g. Aksoy et al. (2019)). In this paper such economic drivers scenarios (i.e. employing the various SSPs) are obtained using the global dynamic economic model MaGE Fouré et al. (2013); Fontagné et al. (2022). This model was developed by J. Fouré, A. Bénassy-Quéré and L. Fontagné in CEPII and is freely available from CEPII’s website [^10] http://www.cepii.fr/cepii/en/bdd_modele/presentation.asp?id=13 in Stata. MaGE assumes that the world consists of economies of individual countries with each country $c$ characterized at time $t$ by a three-factor CES (Constant Elasticity of Substitution) production function with the capital and labour contributions modelled by the Cobb-Douglas parametric form
[…]
where $I$ denotes the GDP, $K$ the capital, $L$ the labour force111111child labour is not modelled by MaGE, $E$ the energy consumption with $c$ denoting the country and $t$ corresponding either to 1-year periods or 5-years periods. The elasticity parameters are assumed to be the same for all countries ($\alpha=0.31$, $\sigma=0.136$) and constant in time, while the parameters $A$ (Total Factor Productivity or TFP) and $B$ (Energy Productivity or Energy Efficiency121212where energy efficiency is defined as an analogous quantity to the ratio between energy consumption and GDP) are assumed to be country specific and temporally varying. The model depends on population primarily through labour force and secondarily through the life cycle savings modelling which is introduced in the modelling of investment. MaGE allows the user to run generalized scenarios concerning the future states of the world economy, compatible with the SSP scenarios as far as population is concerned and produce relevant projections for each SSP scenario. In particular the model included in MaGE allows for the integration of demographics (including education) and economics to allow for predictions concerning population growth and economic quantities such as economic growth rates, GDP, energy consumption, etc.
2.3 A Combined Framework with Climate Scenarios
Since we are investigating food security issues, critical environmental drivers such as temperature, precipitation, water stress, etc. must also be taken into account, given that they affect directly agricultural operations and thus the food system capacity. The so called Representative Concentration Pathways (RCPs) Van Vuuren et al. (2011) determine scenarios regarding the increase of the mean temperature of the planet by the time horizon 2100, taking into account whether certain environmental policies are adopted or not. These scenarios are characterised by the levels of the increase of the mean temperature of the planet and the intensity of the mitigation measures that have to be adopted per scenario (please see Table 7 in Appendix A for the specifications per RCP scenario).
Scenario SSP scenario RCP scenario Environmental Interpretation
SSP1-1.9 SSP1 RCP 1.9 Most optimistic scenario
SSP1-2.6 SSP1 RCP 2.6 Second-best scenario
SSP2-4.5 SSP2 RCP 4.5 Middle of the road scenario
SSP3-7.0 SSP3 RCP 7.0 Baseline of worst-case scenarios
SSP4-6.0 SSP4 RCP 6.0 Best-case of the worst-case scenarios
SSP5-8.5 SSP5 RCP 8.5 Worst-case scenario
Table 3: List of the combined SSP-RCP scenarios investigated in this work
The modern trend in creating realistic scenarios for the future states of the world, is to combine the concept of SSP with that of RCP scenarios. At a first glance, one would naively provide 30 different SSP-RCP scenarios (given the ones illustrated in Table 1 and the standard six RCP scenarios). However, this is not exactly the case since the SSP and RCP scenarios, although referring to different target quantities, are not independent. The Coupled Model Intercomparison Project [^13] https://www.wcrp-climate.org/wgcm-cmip/wgcm-cmip6 (CMIP) has published interesting studies on these climate scenarios by integrating different environmental models and using very dense databases. The combined scenarios (SSP-RCP) that are up to now well tested and available are illustrated in Table 3, consisting of six different scenarios with varying levels of optimism concerning the phenomenon of temperature change Meinshausen et al. (2020). Thinking along the direction of providing food security risk141414in the sense discussed in Section 1 evaluations in a framework which is compatible with both the socioeconomic and climate pathways, we perform estimations according to the six scenarios illustrated in Table 3 building the relevant models for food capacity accordingly (see Section 3).
3 Probabilistic estimations and projections on national food security risk
Predicting future food shortages is of paramount importance to policy makers, as it allows for the adoption of proactive measures, such as for example the restructuring of food production, land use, etc or planning for imports. The design and adoption of effective policies require a deeper understanding of the underlying forces that drive potential major changes in the food system. In this section we propose a (minimal) statistical model, requiring as little as possible and well documented publicly available data, that is able to provide at country level, probabilistic projections of minimum caloric requirements and food system capacity, and hence quantifying food security risk. Population growth, which is the key driver in this model, is treated in a fully probabilistic fashion using the detailed model of Raftery et al. (2013) and its modifications introduced in Section 2. For the modelling task of food security risk, socioeconomic drivers (e.g. GDP, labour, etc), natural resources (e.g. water resources, cropland area, etc) and climate variables (e.g. temperature, precipitation) are involved, however, are not treated in a fully probabilistic fashion. This is partly due to the lack of sufficient computational resources, data or suitable models (such an endeavour would turn the resulting model extremely computationally demanding and also introduce difficulties in its interpretability), so the use of probabilistic scenarios is reserved for the main driver which is population. Clearly, given sufficient resources and appropriate models for the other socioeconomic drivers to build on, these factors can be treated in the same framework. This task is reserved for future research.
Concerning the task of quantifying food security risk, several indicators have been introduced in the literature so far. The UN have proposed the SDG Indicator 2.1.1 or otherwise the Prevalence of Undernourishment (PoU) indicator which is defined [^15] https://unstats.un.org/sdgs/metadata/?Text=&Goal=2&Target as ”an estimate of the proportion of the population whose habitual food consumption is insufficient to provide the dietary energy levels that are required to maintain a normal active and healthy life and is expressed as a percentage”. However, this indicator has been criticised on the grounds that its computation requires quite detailed data, e.g. periodic household surveys should be conducted, detailed information on food acquisition per household and habits, etc. Moreover, to correct potential biases, alternative data on the total amounts of available food for human consumption are needed. As a result, the computation of this indicator is quite a challenging task since rather detailed information is needed at household level, while forecasts concerning the index evolution either in the near future cannot be easily constructed.
Another famous relevant index is the Global Food Security Index (GFSI) [^16] https://impact.economist.com/sustainability/project/food-security-index/. The GFSI considers food affordability, availability, quality and safety, and sustainability and adaptation across 113 countries. The index is a dynamic quantitative and qualitative benchmark model constructed from 68 unique indicators that measure the drivers of food security across both developing and developed countries. Again, the estimation of such an index is far more complex than SDG 2.1.1, since its purpose is to express the state of food security at world (global) scale and unavoidably cannot be projected for prediction purposes.
The modelling approach discussed in the following sections aims at the construction of a reliable and robust with respect to model uncertainty) food security risk index, that can depict if a country is able to cover through its food system the minimum caloric requirements of its population. The index, should rely on data that are easy to obtain and from reliable sources, that are available for many countries over a sufficient depth in time. The proposed index is constructed in a way that can be projected to the future under the combined SSP-RCP scenarios taking into account each scenario’s effect to the underlying quantities that contribute to the indicator value. Note that dietary patterns, food habits and food affordability are not modelled in this version of the indicator.
3.1 Estimating minimum caloric requirements
The need for food is inelastic, in the sense that humans need a minimum and a maximum daily intake of calories for subsistence. The daily calories intake requirements vary per age group, sex, and lifestyle (e.g. level of activity) in a range of 1000 to 3200 calories depending on the category. In Tables 5 and 5 these requirements are shown, as proposed by the HHS/USDA [^17] United States Department of Health and Human Services and Department of Agriculture https://www.dietaryguidelines.gov/sites/default/files/2020-12/Dietary_Guidelines_for_Americans_2020-2025.pdf for the male and female population according to age group and activity level. Given the age structure of the male and female population (i.e. population pyramids per gender) for a country $c$, we may then obtain an estimate for the total daily recommended calories intake at year $t$, through the quantity
(5) $C^{R}_{c,t}=\sum_{a}R_{a}^{f}P_{a,c,t}^{f}+\sum_{a}R_{a}^{m}P_{a,c,t}^{m}$
where $a$ corresponds to the age groups mentioned in Tables 5 and 5, $P_{a,c,t}^{m},P_{a,c,t}^{f}$ denote the total male and female population for the respective age groups and $R_{a}^{m},R_{a}^{f}$ represent the calories requirements given in Tables 5 and 5. This estimate varies, depending on the activity level distribution of the population, however, one may obtain a lower bound for this quantity using the values for $R_{a}^{m},R_{a}^{f}$ for non active individuals or an upper bound using these values for the very active individuals. In this perspective, the quantity $C^{R}_{c,t}$ acts as a proxy for the total daily calorie needs (in terms of a lower or an upper estimate). Clearly, the current estimate may deviate from the actual daily minimum calorie requirements on account of malnutrition issues related to diet patterns, poverty or unequal income distribution, etc. However, as actual data for total calorie needs are not available, we consider $C^{R}_{c,t}$ as a reasonable proxy.
Table 4: Calories Needed Each Day for Boys and Men (Source: HHS/USDA Dietary Guidelines for Americans, 2010)
Age Not Active Somewhat Active Very Active
2–3 years 1,000–1,200 calories 1,000–1,400 calories 1,000–1,400 calories
4–8 years 1,200–1,400 calories 1,400–1,600 calories 1,600–2,000 calories
9–13 years 1,600–2,000 calories 1,800–2,200 calories 2,000–2,600 calories
14–18 years 2,000–2,400 calories 2,400–2,800 calories 2,800–3,200 calories
19–30 years 2,400–2,600 calories 2,600–2,800 calories 3,000 calories
31–50 years 2,200–2,400 calories 2,400–2,600 calories 2,800–3,000 calories
51 years and older 2,000–2,200 calories 2,200–2,400 calories 2,400–2,800 calories
Age Not Active Somewhat Active Very Active
2–3 years 1,000 calories 1,000–1,200 calories 1,000–1,400 calories
4–8 years 1,200–1,400 calories 1,400–1,600 calories 1,400–1,800 calories
9–13 years 1,400–1,600 calories 1,600–2,000 calories 1,800–2,200 calories
14–18 years 1,800 calories 2,000  calories 2,400  calories
19–30 years 1,800–2,000 calories 2,000–2,200 calories 2,400  calories
31–50 years 1,800 calories 2,000  calories 2,200  calories
51 years and older 1,600 calories 1,800  calories 2,000–2,200 calories
Table 4: Calories Needed Each Day for Boys and Men (Source: HHS/USDA Dietary Guidelines for Americans, 2010)
Table 5: Calories Needed Each Day for Girls and Women (Source: HHS/USDA Dietary Guidelines for Americans, 2010)
The population model (see Section 2.1) provides accurate probabilistic predictions for the population pyramid, i.e., for the quantities $P_{a,f,c,t},P_{a,c,m,t}$. Using this model, a number (let us say $M$) of different realizations for the population pyramid in terms of data batches of trajectories are obtained, concerning the evolution of both female and male population per age group over the time period $[T_{0},T]$, i.e.
$P_{f}(\tau,\mathcal{J}):=\left\{P_{a,c,f,t}^{(j)}\,:\,j\in\mathcal{J},\,\,t\in\tau\right\},\,\,P_{m}(\tau,\mathcal{J}):=\left\{P_{a,c,m,t}^{(j)}\,:\,j\in\mathcal{J},\,\,t\in\tau\right\}$
where $\tau=T_{0},...,T$ and $\mathcal{J}=\{1,2,...,M\}$. As already stated, the uncertainty effects are properly accounted for in these trajectories and in accordance to past data. Taking a slice of, e.g., $P_{f}(\tau,\mathcal{J})$ at a fixed time $t^{\prime}\in[T_{0},T]$, will provide the sample $P_{f}(t^{\prime},\mathcal{J})$ consisting of $M$ possible realizations of the random variable $P_{a,c,f,t^{\prime}}$ which can provide useful information concerning its distribution (i.e. moments, quantiles, etc). In fact, the general trajectories can be classified according to various criteria that characterize the SSP scenarios (see Section 2.2) so as to obtain subsets of the trajectories which are compatible with the various SSP scenarios. Using the trajectories for each scenario we may obtain conditional means or quantiles for the conditional distribution of the quantities $P_{a,c,f,t^{\prime}},P_{a,c,m,t^{\prime}}$ per SSP scenario. This procedure allows us to have a detailed probabilistic scenario-based description of the possible evolution of future population related quantities.
After obtaining the trajectories and probabilistic scenarios for the population and its age structure, using the estimate (5) for the minimum caloric requirements, we may generate similar probabilistic scenarios for $C^{R}_{c,t}$ for the future. To this aim, we have to use the generated data batches $P_{f}(\tau,\mathcal{J})$, $P_{m}(\tau,\mathcal{J})$, and feed them into formula (5) to generate trajectories $C(\tau,\mathcal{J}):=\{C^{R}_{c,t,j}\,:\,t\in\tau,\,\,j\in\mathcal{J}\}$ where $C^{R}_{c,t,j}$ denotes the calorie requirements for the country $c$ at time $t$ according to the population pyramid generated in the $j$-th trajectory of the sample. These data batches will be subsequently used to generate samples for projections of $C^{R}_{c}$ on various future dates $T_{0}\leq t^{\prime}\leq T$, and from those as described above, probabilistic information on this important quantity will be generated. Clearly, when this quantity is needed in the context of SSP scenarios, the relevant trajectories for the population quantities corresponding to these scenarios will be employed to the generation of trajectories in (5) for estimating the minimum caloric requirements.
Remark 3.1.
Minimum calories intake is modelled by the nominal value of the calories that are required by the population to survive, based only on the population structure and the nominal values for calories estimated by nutrition experts. Neither diet patterns or food product substitutions (e.g. when pasta is unavailable, replacing with rice, etc), or access to food are taken into account in this approach. The resulting estimate $C^{R}_{c,t}$ stated in (5) does not necessarily coincide either with the minimum nutritional needs or the food consumption of the country under study. Nutritional needs are subject to dietary patterns, available crops or food types. The estimation of minimum nutritional needs is a much more complex task than the estimation of the minimum calorie intake, since it relies on several aspects of food demand. Similarly, food consumption also involves aspects of food demand and supply (see e.g. Hasegawa et al. (2015); Valin et al. (2014); Van Dijk et al. (2021)). However, the estimate $C_{c,t}^{R}$ provided here, depends directly only on population structure of the country under study and does not take into account other factors, in an attempt to provide a rough, but as reliable as possible, estimate for the caloric requirements of the country’s population for subsistence. Clearly, a more detailed model could be considered as a next step in this modelling component, employing more detailed data concerning the nutrition patterns and habits, and the effects of active policies for increasing food affordability of the general population.
3.2 Modelling food system capacity
Minimum food requirements is only the one side of the food security risk discussion, with the food system capacity (as a proxy for food supply) being its indispensable counterpart. Potential food security issues may arise when the food system capacity (in terms of calories available for human consumption) cannot cover the minimum calorie requirements of population, $C^{R}_{c,t}$, for a particular country at a certain point in time. Clearly, socioeconomic conditions, natural resources availability and of course climate conditions (directly or indirectly) affect the capacities of the national food systems. Other aspects of food security, such as physical, social and economic access to food, sustainability and nutrition, although extremely important Burlingame (2014); Garnett (2013), are bound to be subsequent considerations and at the same time policy arenas, where food security issues arising from deficient food system capacity can be mitigated or reversed. Our approach for estimating food security risk based on long-term drivers is complementary to previous attempts, that are usually focused on short-term drivers of food security (see e.g. WHO et al. (2021)). Following Ericksen’s food system framework Ericksen (2008), the main long-term drivers of food system outcomes (here calories available for human consumption) are grouped into global environmental change drivers (e.g. changes in land cover, climate variability, water and nutrient availability, etc.) and socioeconomic drivers (e.g. changes in demographics, economics, technology, etc.). The High Level Panel of Experts on Food Security and Nutrition further classifies these drivers into six categories: biophysical and environmental; technology and innovation; economic and market; political and institutional; socio-cultural; and demographic Fanzo et al. (2017).
Based on the above, to examine the effect of SSP-RCP scenarios on food security, we opt for a minimal, but reasonable and identifiable, model with a generic structure, where food system capacity is subject to the effects of the aforementioned drivers. In particular, we consider changes in socioeconomic (i.e. population, income, labour) and environmental (i.e. land, water resources, precipitation, temperature) factors that affect the food system capacity mechanism in the long-run, either directly or indirectly, through the food system activities (production, consumption, markets, trade, etc.). The above drivers are the ones for which data on future projections, consistent with each SSP-RCP scenario, are readily available or can be derived. Of course, other relevant economic drivers such as land use, labour force occupied in the food production sector, etc. could also be included Gaitán-Cremaschi et al. (2019); Fanzo et al. (2017); Van Berkum et al. (2018), but such detailed data are not usually available, especially for developing (non-OECD) countries. Also, even if available, predicting future values of such quantities under the various scenarios, and in a form compatible with the probabilistic scenarios framework, may be difficult or computationally impossible, at least with the current state of the art of economic modelling and computational resources, respectively.
To better incorporate the aforementioned features, the food system capacity is modelled by two terms corresponding to distinct contributions: (a) the domestic food production and (b) the food demand effect, which is introduced indirectly to our model through the imported and exported quantity of food. Domestic food production quantity corresponds to the total quantity of food produced within country $c$, which crucially depends on the natural resources of the country, such as land and water resources, as well as the labour force occupied in the food production sector (agriculture or livestock) and the states of the economy. On the other hand, the aspects of food demand represented by food quantity imports and exports, rely mostly on socioeconomic factors like GDP or the population level. In this perspective, the food system capacity is estimated as the aggregation of these two major contributions where other relevant quantities are determined on a bottom-up, two layer modelling approach. Historical data for food system capacity at country level are obtained by the FAO database [^18] FAO STAT Web Database: https://www.fao.org/faostat/en/#data/SCL (referred to as total food supply in the database).
In this spirit, an identifiable modelling approach for the food system capacity is represented by the two-layer scheme:
(10) Upper Layer $\mbox{\bf Upper Layer }\left\{\begin{array}[]{ll}Q^{FSC}_{c,t}&=f_{1}\left(t,Q^{Dom}_{c,t},Q^{Exp}_{c,t},Q^{Imp}_{c,t}\right)\\ Q^{Dom}_{c,t}&=f_{2}(t,P_{c,t},I_{c,t-1},L^{Agr}_{c,t},W_{c,t},A_{c,t},T_{c,t})\\ W_{c,t}&=f_{3}\left(t,P_{c,t},I_{c,t-1},Q^{Dom}_{c,t},A_{c,t},T_{c,t},Pr_{c,t}\right)\\ A_{c,t}&=f_{4}(t,P_{c,t},I_{c,t-1},Q^{Dom}_{c,t-1})\end{array}\right.$
(14) Lower Layer $\mbox{\bf Lower Layer }\left\{\begin{array}[]{ll}Q^{Imp}_{c,t}&=g_{1}(t,P_{c,t},I_{c,t-1})\\ Q^{Exp}_{c,t}&=g_{2}(t,P_{c,t},I_{c,t-1})\\ L^{Agr}_{c,t}&=g_{3}(t,P_{c,t},I_{c,t-1},L_{c,t})\end{array}\right.$
where the upper layer concerns the modelling of quantities:
  • $Q^{FSC}_{c}$: the food system capacity of country $c$ at time $t$,
  • $Q_{c}^{Dom}$: the domestically produced food quantity,
  • $W_{c}$: the level of water stress of country $c$ at time $t$, i.e. the freshwater withdrawal in percentage of the available freshwater resources (according to the SDG Indicator 6.4.2191919https://www.fao.org/publications/card/en/c/CA8358EN/),
  • $A_{c}$: the land area occupied for agricultural activities,
while the lower layer concerns the modelling of the quantities:
  • $Q^{Exp}_{c},Q^{Imp}_{c}$: the exported and imported food quantities, respectively, and
  • $L_{c}^{Agr}$: the labour force occupied in the food production (agricultural) sector,
which are subsequently fed into the upper layer to produce food system capacity estimates.
The splitting of the models in two layers concerns the better presentation of the total modelling approach for food system capacity and does not introduce extra assumptions and restrictions concerning the estimation procedure. The lower layer quantities depend mainly on socio-economic drivers and potential projections for these quantities are more dependent on the various SSP scenarios. On the other hand, the upper layer quantities depend both on general socioeconomic and climate drivers, therefore potential projections on these quantities depend both on SSP and RCP scenarios. More specifically, all quantities modelled in the lower layer depend directly on population ($P$), labour ($L$) and GDP ($I$), whose evolution differs depending on the SSP scenario that is considered for the future (SSP-dependent quantities). On the other hand, the quantities modelled in the upper layer depend directly on both population ($P$) and GDP ($I$), but also on environment related quantities, such as temperature ($T$), precipitation ($Pr$) etc. whose future evolution depends on the RCP scenarios. Moreover, no interdependencies are considered between the various factors in the lower layer, while on the upper layer interdepencies between the modelled quantities are considered (for a graphical illustration of the modelling approach please see Figure 3). Extra assumptions (e.g. specific forms for evolution on agricultural labour force taking into account technological indicators) or constraints (e.g. policy or physical restrictions on using land and water resources) on the relations between the modelled quantities may be introduced by appropriate choices for the functions $f_{1},f_{2},f_{3}$ and $g_{1},g_{2},...,g_{4}$. Note that although for the case studies considered in this paper (please see Section 4) only Cobb-Douglas type models are used, these models could be substituted with any other econometric modelling approach (e.g. Translog models, CES production functions, etc) depending on the data availability.
Figure 3: The modelling approach concerning the food system capacity and minimum calorie requirements, illustrating inter-dependencies between the various model components and the main effects introduced by the various socioeconomic and climate drivers
The modelling scheme (10)-(14) attempts to describe the mechanism resulting to the observed food system capacity relying on a set of socioeconomic quantities and environmental drivers. Population affects the system through various routes; it clearly affects the labour force as well as the income. Natural resources as well as environmental resources affect the domestic food production quantity. The generated food capacity for the system, allows for both domestic and imported food supply, and this depends on the level of the food quantity generated by the domestic mechanism, the effect of food demand which is captured by the food imports and exports terms, while both depend on economic quantities - GDP - which may play an important role on the ability of the country to acquire imported goods. The modelling task of these terms is performed in a regional scale, not taking into account global effects (e.g. interdependencies and interactions of the world market) which could be a direction for future work. Note also that GDP is included as a single-step delay to capture its effect on food system activities (e.g. domestic production, and related food imports and exports) but at the same time avoid reverse causal effects of these activities on GDP (Zestos and Tao (2002)). Our choice of model implicitly assumes that the drivers not included in it remain constant, while food system activities (e.g demand patterns, production technology, crop usage, food trade balance, etc.) are not disrupted (in the sense of a structural break) by the included drivers (changes in population, GDP and environmental variables, according to the SSP-RCP scenarios). As explained in the following section, under these assumptions, we are able to make predictions concerning food capacity, which will be necessary to establish a risk measure that can provide information to policy-makers, in advance for changes that need to be made in order to reduce food security risks.
3.3 A food security risk index within and across socio-economic scenarios
Based on the modelling approach presented in Sections 3.1 and 3.2, for the minimum caloric requirements ($C^{R}$) and the food system capacity ($Q^{FSC}$), we introduce in this section a new food security risk index. Combining these models we attempt to assess the food security risk at national level adopting an approach that allows for estimating/evaluating the future food security risk under different SSP-RCP scenarios, while requiring the input of less detailed data, as compared to other well established indices (such as the SDG 2.1,1 Indicator or the GFSI). The proposed indicator treats robustly the uncertainty inherent in the estimation of the key factors which affect the food security phenomenon (i.e. future population estimates which are provided in terms of probabilistic projections) either within or across the scenarios that are considered.
We take this opportunity here to clarify what we mean by within and across scenarios. For various reasons (e.g. the long term horizons involved in the modelling of the phenomenon, or lack of sufficient data etc) it is not always possible for the decision maker to be aware of the exact scenario that will be materialised. Moreover, due to the long time horizon and the dynamic nature of the phenomenon, it may be that we start within the range of one scenario and in the course of time, as the phenomenon evolves, we enter within the range of a different scenario. This means that we should be able to make predictions concerning food security not only within a particular scenario (an SSP or an SSP-RCP scenario, i.e. referring to the “within scenario” case) but also taking in to account the possibility of possible multiple scenarios, or even transitions from a scenario to another. We will refer to the latter situation using the terminology “across scenarios”, i.e. the situation where SSP or SSP-RCP mixed scenarios are considered, where each scenario in the set is considered as probable with a specified probability. The rationale behind this approach, is to provide a risk assessment tool capable of supporting the decision making process of policy makers, developing a working framework that allows to take into account the effects of important drivers and the uncertainty propagated to the food security risk assessment task.
Keeping in mind the aforementioned considerations we define the new food security risk index.
Definition 3.2 (FSRI).
Given the minimum caloric requirements ($C^{R}$), the food system capacity ($Q^{FSC}$) and the water stress ($W$) of a country $c$ for a specific year $t$, the Food Security Risk Index (FSRI) is defined as
(15) $I^{FS}_{c,t}:=\left[\frac{1}{1+\gamma}\left(\frac{C^{R}_{c,t}}{Q^{FSC}_{c,t}}\right)+\frac{\gamma}{1+\gamma}W_{c,t}\right]\cdot 100\%$
where the sensitivity parameter $\gamma\in[0,\infty)$ denotes the relevant cost for country $c$ in acquiring extra water resources.
The proposed index expresses the percentage ratio between the minimum caloric requirements to the food system capacity capabilities at national (or regional) level, weighted by the level of water stress with respect to the sensitivity parameter $\gamma$. In the special case where the water stress risk is not taken into account (i.e. $\gamma\to 0$, assuming extremely low cost in acquiring extra water resources besides the available ones from the country, or a case of a country that will never face water stress issues), high capacity of the food system $Q^{FSC}$ comparing to $C^{R}$ leads to lower indicator values, while low $Q^{FSC}$ values comparing to $C^{R}$ should lead to increased food security risk. However, water stress risk in general significantly affects food security since most food production activities relies on water resources management and is of crucial importance, especially in areas where water resources are very limited, e.g. North Africa region. Clearly, the determination of the relevant sensitivity parameter $\gamma$ should be done with special care by the policy maker, taking into account the current and the future status of the under study area concerning the pressure on water resources and their necessity to the national or regional food production activities.
Involving water stress explicitly in the index, although is already accounted for implicitly in the other quantities $C^{R}$ and $Q^{FSC}$, contributes in its robustification with respect to potential model misspecification issues that may arise. Unavoidably, the modelling tasks of water stress and food system capacity are subject to model misspecification errors that may occur due to sudden changes in the socio-economic conditions in an area (e.g. political arrests, wars), pandemics (e.g. COVID), climate change effects (e.g. natural hazards), etc that may lead to temporally misleading estimates (overestimation or underestimation of various drivers’ effects) concerning the food demand and food supply patterns. Subsequently, poor fitting results due to such reasons, may provide unrealistic projections, i.e. enormous increase in food production quantities which may not be supported either by the capacity of the food production mechanism or the availability of natural resources (e.g. water resources). The explicit inclusion of the water stress term aims to partly treat this matter by directly introducing the risks related to the water resources availability, especially in the case where the country under study is already water stressed202020any country with water-stress indicator value above 40% is considered as highly water-stressed. According to World Resources Institute (WRI) [^21] https://www.wri.org/insights/highest-water-stressed-countries several levels for water-stress are defined with respect to the interval that the water stress indicator lies (i.e. the term $W$ in our index). In particular, the following characterizations are introduced: extremely high water stress for values greater than $80\%$, high water stress in the interval $40\%-80\%$, medium-high water stress in the interval $20\%-40\%$, low-medium water stress in the interval $10\%-20\%$ and low water stress for values less than $10\%$. The countries under study in this work, Egypt and Ethiopia, display two different patterns concerning pressure on water resources, at least for the period 2001-2018 (Figure 4). Egypt is considered as extremely high water stressed country according to the WRI definitions ($W>80\%$), and about the last two decades water stress oscillates around $120-125\%$. On the other hand, Ethiopia at early 2000s was considered a low-medium water stressed country, however it presents a clearly increasing trend and at the year 2018, the water stress is about $35\%$ characterizing the country as a medium-high water stressed one. Therefore, it seems that the pressure on water resources could provide quite different pictures even in neighbouring countries, and the inclusion of this information to the food security risk index might help in representing more accurately the actual situation.
Figure 4: Level of water stress (SDG 6.4.2 Indicator) for Egypt and Ethiopia for the time period 2001-2018.
As our proposed index treats the risks in a more aggregate fashion than other detailed indices (such as SDG 2.1.1) careful and meaningful choice of the parameter $\gamma$ is required if we expect our simple index to reproduce features of other more complex indices. We must emphasise here that the simple index proposed here is suitable for projections into the future, unlike the other more complex indices. Hence, tuning the sensitivity parameter $\gamma$ in such a way such that the proposed index reproduces for past data features showing in the more complex indices, we enhance our faith in the predictive ability of our index. Given the WRI specifications related to the water stress intensity, we may provide different perspectives concerning the sensitivity of the proposed food security risk index concerning the pressure on water resources. A quite rational rule would be one to consider that the sensitivity parameter value evolves linearly (or even more steepier) with respect to the distance of the previously recorded water stress value from a safety threshold. So, if previously recorded water stress is below the safety threshold the water stress component is deactivated in the index ($\gamma=0$), otherwise $\gamma$ represents the distance from the safety threshold. For illustration purposes, we consider three different perspectives: (a) the very conservative perspective (VC) that penalizes when water stress surpasses the low-medium level ($>10\%$), (b) the less conservative perspective (LC) when the medium level ($>20\%$) is violated and (c) the no conservative perspective (NC), where only pressure beyond the high water stress threshold ($>40\%$) is taken into account and activates the relevant water risk term in the index. These perspectives allows for the determination of the sensitivity parameter in a dynamic fashion, relying on the historical water stress evolution. The perspectives introduced above are represented by the following rules:
(16) $\left\{\begin{array}[]{ll}\mbox{\emph{Very conservative perspective (VC):}}&\gamma_{c,t}:=\max(0,W_{c,t-1}-0.10)\\ \mbox{\emph{Less conservative perspective (LC):}}&\gamma_{c,t}:=\max(0,W_{c,t-1}-0.20)\\ \mbox{\emph{No conservative perspective (NC):}}&\gamma_{c,t}:=\max(0,W_{c,t-1}-0.40)\end{array}\right.$
The obtained values for $\gamma$ under these three perspectives are illustrated in Figure 5 for Egypt and Ethiopia for the time period 2001-2018. Since Egypt is a extremely high water stressed country, all three rules lead to strictly positive values for $\gamma$ of that evolve on exactly the same manner with different horizontal position (translated lower or higher depending on the perspective). On the other hand, for Ethiopia the NC perspective does not allocate any positive value on $\gamma$ since the water stress index for this period is below the activation threshold of $40\%$. The LC perspective provides positive values for $\gamma$ after the year 2004 with increasing trend, while under the VC perspective positive values are allocated on $\gamma$ for the whole period with also increasing trend.
Figure 5: Evolution patterns for the sensitivity parameter $\gamma$ according to the perspectives stated in (16) for Egypt and Ethiopia for the time period 2001-2018.
Figure 6: Illustration of the SDG 2.1.1 (PoU) index and FSRI for Egypt (upper panel) and Ethiopia (lower panel) for the period 2001 - 2018 under the perspectives stated in (16).
Although even more sophisticated approaches can be considered in the method that the two risks are aggregated, it seems that this quite simple threshold-based approach, allows to qualitatively approximate well established indices like PoU (SDG 2.1.1) that require more detailed data. This fact is illustrated in Figure 6 where the PoU indicator and the proposed index (FSRI) for different values of $\gamma$ are depicted for the both countries under study. Note that for comparison purposes, the case where water stress is not taken into account to the food security risk calculation ($\gamma=0$) is also included. Clearly, for Egypt the tendency of the PoU indicator is better represented by FSRI for any of the choices that allocate positive values on $\gamma$ comparing to the case where the term of water risk is excluded from the calculation ($\gamma=0$). Moreover, for Ethiopia the nice behaviour of the FSRI in approximating the tendency of PoU is also observed without any major misspecifications for any choice of $\gamma$. However, this is expected to change if the country enters the highly water-stressed stage, where the water risk component of the index is expected to play more important role in the better representation of the food security risk situation.
Following the SSP-RCP scenarios framework introduced in Section 2 combined with the modelling approaches discussed in Sections 3.1 and 3.2, we are able to provide/generate probabilistic scenarios for the food security risk index $I^{FS}$ as follows:
  1. 1.
    Scenarios for Minimum Caloric Requirements
    We use the procedure in Section 3.1 to provide future estimates and SSP compatible scenarios for the minimum calorie requirements at national level $C_{c,t}^{R}$ based on the detailed population evolution scenarios.
  2. 2.
    Scenarios for Food System Capacity
    Model layers 10-14 are calibrated using available historical data from a sufficiently large period for the country $c$ to estimate the profile of country $c$ (i.e. obtaining the relevant country-specific parameters). (b) The fitted model is then applied for predicting the future food system capacity $Q^{FSC}_{c,t}$ using the probabilistic scenarios (and the relevant trajectories) for the population (see Section 2.2) along with projections for the future GDP per capita as obtained from the global macroeconomic model MaGE Fouré et al. (2013); Fontagné et al. (2022).
  3. 3.
    Scenarios for the Food Security Risk Index
    Using the trajectories for $Q^{FSC}_{c,t}$, $C_{c,t}^{R}$ and $W_{c,t}$ obtained in the previous steps we construct trajectories compatible with the various SSP-RCP scenarios for the index $I_{c,t}^{FS}$ and use the trajectories to provide statistical information for the index in the various scenarios.
Recalling our modelling framework (please see Figure 3), we realize that the generated values or trajectories for the food security risk indicator depend on the set of key factors
$Z_{c,t}:=(P^{f}_{a,c,t},P^{m}_{a,c,t,},T_{c,t},Pr_{c,t})$
. These are the main stochastic factors which introduce uncertainty to the modelled quantities that lead to the estimation of $I^{FS}_{c,t}$. This relation is represented through a risk mapping $Z\mapsto I^{FS}(Z)$, connecting the random risk factors collected in $Z$ with the risk output $I^{SF}$, through the models presented in Sections 3.1 and 3.2 for a country $c$, i.e. expressing $I^{FS}$ in terms of $I^{FS}:=\Phi_{c}(Z)$. Let us denote by $Q$ the probability measure under which the sample of trajectories for $Z$ is generated (i.e. coinciding with one of the SSP-RCP scenarios in Table 3). From now on we will identify the output of the various scenarios, with respect to the risk factors $Z$, with probability measures $Q$, or equivalently with probability distributions for the risk factors, which in turn will induce via the risk mapping $\Phi_{c}$ probability measures or probability distributions for the food security index $I^{FS}$. Employing the standard framework of risk management, we may use the above setting in defining a risk measure associated with food security, as quantified by the index $I^{FS}$. In the case where a probabilistic model $Q$ for the description of the risk factors $Z$ is universally accepted, then, the best estimate for the risk would simply be the expectation of the risk mapping under the probability model $Q$, i.e. ${\mathbb{E}}_{Q}[I^{FS}]={\mathbb{E}}_{Q}[\Phi_{c}(Z)]$. In the present context, this would correspond to choosing one particular SSP-RCP scenario, the most plausible one, which would be identified with a probability measure $Q$ for the risk factors $Z$, and then using the risk mapping obtain an estimation of the risk measure in terms of ${\cal I}^{FS}={\mathbb{E}}[\Phi_{c}(Z)]$. Having more than one scenario, would correspond to obtaining a set of possible probability measures for the evolution of the risk factors $Z$,
(17) $\mathcal{Q}$ =$=\{Q_{1},Q_{2},Q_{3},Q_{4},Q_{5},Q_{6}\}$
=:$=:\{Q_{SSP_{1}-1.9},Q_{SSP_{1}-2.6},,Q_{SSP_{2}-4.5},Q_{SSP_{3}-7.0},Q_{SSP_{4}-6.0},Q_{SSP_{5}-8.5}\},$
and upon selecting any of these, the food risk security index is calculated as
(18) $\mathcal{I}^{FS}_{c,t}=\mathbb{E}_{Q}\left[I^{FS}_{c,t}\right]=\mathbb{E}_{Q}\left[\Phi_{c}(Z_{c,t})\right].$
The setting above allows for the food security risk estimation within each one of the considered socioeconomic-climate probabilistic scenarios, with the term within meaning that we base our estimations on a single scenario out of the set ${\mathcal{Q}}$. However, it is not always feasible to identify with certitude the exact scenario that we are experiencing, or (as happens in the case of modelling within long time horizons) to know a priori that we will remain within one scenario due to the dynamic nature of the phenomenon. In such cases, one may wish to blend the various future perspectives in order to provide a risk assessment concerning the food security issue across scenarios. This is more realistic since in practice, he existence of a single universally accepted probability model (scenario) $Q$ is unusual. Instead of this, the set of probability models under different possible scenarios, ${\cal Q}$, provides different aspects of information concerning the evolution of the risk factors $Z$. This brings us to the realm of Knightian uncertainty Knight (1921), in which their is not a unique, universally acceptable probability measure for the description of the evolution of the risk factors affecting the phenomenon under study. This is a situation which requires a better way of providing the best estimate for the risk, than simply using the expectation ${\mathbb{E}}_{Q}[\Phi_{c}(Z)]$ under a unique probability measure. This would constitute what we call an estimator for the risk across scenarios; i.e. a measure of the risk of food security, under the uncertainty as to which scenario (equiv. model) will actually materialize. A proposal along these lines which is currently well accepted by the community of risk management, is within the context of convex risk measures and their robust representation (see e.g. Detlefsen and Scandolo (2005); Föllmer and Schied (2002); Frittelli and Gianin (2002)). Within this framework we calculate the risk through the variational representation
(19) $\mathcal{I}^{FS}_{c,t}:=\rho(I^{FS}_{c,t})=\sup_{Q\in\mathcal{P}}\left\{{\mathbb{E}}_{Q}[\Phi_{c}(Z_{c,t})]-a(Q)\right\},$
[…]
Having adopted the fundamental conceptual framework of treating each scenario as a different probabilistic model (probability measure) for the risk factors $Z$ (mainly population factors in this study) we may now answer the question of robust estimation of the quantity of interest $I^{FS}_{c,t}$, for fixed $t$, using the convex risk measure of the form stated in (19). Following the suggestion in Papayiannis and Yannacopoulos (2018) and Petracou et al. (2022), in order to distinguish between the variou scenarios we adopt a metric space setting, according to which we consider the various scenarios in their natural setting, i.e., the metric space of probability measures, with a suitable metric $d(\cdot,\cdot)$, chosen so as to differentiate between the various probability measures in ${\cal Q}$, associated with different scenarios.
The risk measure is chosen so as to display uncertainty aversion, a property that is guaranteed (see Petracou et al. (2022), see also Papayiannis and Yannacopoulos (2018)) by choosing the penalty function
(20) $a(Q)=\frac{\theta}{2}\sum_{i=1}^{6}w_{i}d^{2}(Q,Q_{i}),$
where $Q_{i}$ are the probability measures (probabilistic scenarios) included in the set $\mathcal{Q}$ stated in (17), where $w_{i}$, $i=1,2,\ldots,6$ are (credibility) weights associated to each scenario (these can be subjective and associated to expert opinion or objective i.e. derived from evidence from the data and possibly updated through a suitable learning scheme), $d(\cdot,\cdot)$ is a metric in the space of probability measures and $\theta>0$ is the uncertainty aversion parameter, modelling the propensity of the decision maker to deviate from the probability models in ${\cal Q}$. A suitable choice for $d$ is the Wasserstein metric (see e.g. Santambrogio (2015); Villani (2021)) which is directly related to the misspecification error of the random variable $I^{FS}$ if a different probability model for $Z$ is chosen in place of the true model. Moreover, the choice (19), with (20) combined with the Wasserstein metric, allows for efficient numerical calculation of the risk measure ${\cal I}^{FS}$ for a wide class of probability models. For the interesting case where all plausible models are included in $\mathcal{Q}$ (corresponding to the limit $\theta\to\infty$) we obtain the approximation of ${\cal I}^{FS}$ by
(21) ${\cal I}^{FS}_{c,t}=\mathbb{E}_{Q^{*}}[I^{FS}_{c,t}]={\mathbb{E}}_{Q^{*}}[\Phi_{c}(Z)],$
where $Q^{*}$ is the barycentric probability model over all scenarios in $\mathcal{Q}$ with respect to the weights $w$ (under the Wasserstein distance sense, see e.g. Agueh and Carlier (2011)). The risk measure ${\cal I}^{FS}$ as stated above, provides a robust estimate for the food security risk, across scenarios, in the limit of deep uncertainty. This is the main approximation we will be using in this work, however further approximations are possible, if required, using the $\Delta$-approximation of the risk mapping $Z\to\Phi_{c}(Z)$ (for details please see Papayiannis and Yannacopoulos (2018)).
The algorithmic approach in estimating the food security risk index across scenarios can be thus summarized as follows:
Algorithm 3.3 (Food Security Risk Estimation Across SSP-RCP scenarios).
  1. 1.
    Fix a certain country $c$ and a time $t$.
  2. 2.
    Define the risk factors $Z=(P_{a,f},P_{a,m},T,Pr)$ and using the procedure described in Section 2.2 obtain probabilistic scenarios for $Z$ and determine the corresponding probability models $Q_{i}$, $i=1,\ldots,6$ through the generated samples.
  3. 3.
    By estimating the model layers described in (10)-(14) and using the expression (5) obtain the risk mapping $Z\mapsto I^{FS}_{c}=\Phi_{c}(Z)$ for $I^{FS}_{c}$ as defined in (15).
  4. 4.
    Obtain the barycentric scenario $Q^{*}$ provided a choice of weights $w$ allocating the degree of realism of the policy maker to each one of the scenarios under consideration.
  5. 5.
    Using Monte-Carlo simulation estimate ${\cal I}_{c}^{FS}={\mathbb{E}}_{Q^{*}}[\Phi_{c}(Z)]$.
4 Food security risk assessment for Egypt and Ethiopia under the SSP-RCP framework
4.1 Data and model assumptions
We follow the modeling approach presented in Section 3 for describing (a) the minimum caloric requirements (Section 3.1) and (b) food system capacity (Section 3.2) in order to provide estimates concerning food security risk for the cases of Egypt and Ethiopia through the risk indicator introduced in Section 3.3. For the estimation of the minimum calories intake component, it suffices to use the probabilistic estimates concerning the population structure evolution for Egypt and Ethiopia, since this quantity depends only on population (see the relevant discussion in Section 3.1). On the other hand, food system capacity at country level is described in terms of the two-layer model stated in 10-14, therefore relevant training data for the model calibration for both countries are required. For the model calibration task, historical data from the period 1990-2018 were used, since for this time period data are available or all quantities of interest that are taken into account in the modelling procedure. The required data for the model fitting procedure for Egypt and Ethiopia were mainly collected by the open database available from FAO, which includes historical data for food system capacity, food quantity imports and exports and land area used for agricultural purposes at national scale. Climate data and in particular, average temperature and average precipitation at national scale were obtained by the World Bank’s Climate database [^23] https://climateknowledgeportal.worldbank.org/ while water-stress records/estimations were obtained from the World Bank Database [^24] https://data.worldbank.org/. The variables that were used for the model calibration task and the sources from which the relevant data were acquired are illustrated in Table 8 in the Appendix.
Concerning the model fitting procedure, Cobb-Douglas type parametric models were employed for the estimation of all the required parameters. For a full model description please see Section LABEL:app-4 in the Appendix. The fitting/estimation procedure has been performed separately for each country (not as panel data) using Ridge-type penalization to the typical OLS estimation approach, to treat potential instabilities caused by collinearity effects of the predictors. Note also that models are fitted separately to data and they are not fitted as a system of equations since different set of effects are considered for each quantity. The obtained model parameters for Egypt and Ethiopia are illustrated in Tables 9 and 10 in the Appendix with the relevant goodness-of-fit results. Concerning the food system capacity profiles, significant qualitative differences are not observed between the two countries. The technological and the year trend effects are at the same levels as far as the domestic food production is concerned. The only qualitative difference that is observed, concerns the effect of food exports where for Egypt the relevant coefficient displays a negative value, which possibly represents the difference in the food production capabilities and patterns between the two countries. Exported and imported quantity models seems to be quite similar for Egypt and Ethiopia. The food production mechanism of Egypt seems to be negatively related to potential temperature increases. Moreover, a negative dependence between the labour force occupied in food production activities in Ethiopia and GDP is observed, displaying potential social trends when income increases. However, most of the effects should be considered as aggregate effects and it is an interesting question for future research to see if these effects persist in finer scale models, e.g. involving a differentiation between food products or regions. However, such a task is subject to the availability of more detailed data, which are not currently available.
4.2 Food security risk estimation within SSP-RCPs
Following the modelling of minimum food requirements and food system capacity, estimates for both quantities concerning the time period 2019-2050 for Egypt and Ethiopia under each SSP-RCP scenario are produced (for the list of scenarios presented in Table 3). Population projections under each SSP scenario are directly available through the probabilistic model presented in Section 2.2. Concerning the other socioeconomic indicators, there exist several macroeconomic models that, based on population projections, provide projections for various socioeconomic indicators of interest under the SSP scenarios. In this paper, the MaGE model (see Section 2.2 and Fouré et al. (2013); Fontagné et al. (2022) for more details) is employed for this task. The MaGE model based on the UN and IIASA [^25] https://iiasa.ac.at/ databases provides projections up to year 2100 for the socioeconomic activities of all countries of the world (being a global model) under each one of the SSP scenarios. In this paper, the MaGE model is extended in order to provide projections of the socioeconomic variables that are compatible with the probabilistic population projections. This is achieved by substituting the standard population inputs in MaGE with the conditional means for population under each SSP, obtained in terms of the simulated trajectories (accordingly assigned to the appropriate SSP scenario using the methodology introduced in Section 2.2), to provide estimates for GDP and labour, compatible with the philosophy of SSP probabilistic modelling. Note that this task could be performed trajectory-wise from the population scenario database i.e., generating individual trajectories using MaGE for the socioeconomic quantities in question, one for each individual trajectory for the population evolution, within scenarios. This would lead to a sample of trajectories per scenario for the socioeconomic quantities, which could then be used for the generation of probabilistic projection for the socioeconomic quantities, in a manner similar to that described in Sections 2 and 3. However this would be quite expensive in computational time and possibly colliding with the general modelling philosophy of MaGE which is based on pointwise estimates, therefore it would essentially require a brand new macroeconomic model.
For the projection tasks of this work, these two worlds are combined using the future estimates for GDP evolution as provided by MaGE under each one of the SSP scenarios, incorporating the conditional means for the population related quantities as obtained by the full probabilistic population model, keeping in mind the possible limitations and drawbacks from this approach. Climate drivers and in particular temperature and precipitation (in annual basis), are provided for each one of the scenarios in Table 3 from the Climate Change Knowledge Portal online database. Projections about the land area that will be occupied for agricultural activities, are performed under the assumption that no changes to the current land use policies are made. However, constraints as to the maximum area occupied for agricultural purposes according to each country capabilities in providing agricultural land, are applied. In this direction, for Egypt and Ethiopia we assign a maximum land area value that can be allocated to agriculture until 2050 with respect to the current capabilities of the countries. Therefore, to determine each country’s profile in agricultural land use, we estimated its trend for the last two decades, restricted by the upper bound that has been set by the physical limitations in land use. In this way, since land use depends on population, economy and climate factors (see model (14)), different estimates are obtained under each SSP-RCP scenario. Although more sophisticated models can be considered, taking into account thinner data about land use, under which the effect of land policies might be assessed by the proposed approach, we believe that such a task is beyond the scopes of this work. The key socio-economic and climate drivers which where used for the projections and the related models and sources that they were obtained from, are illustrated in Table 11 in Appendix C while elements concerning the uncertainty quantification of all main drivers that contribute to our modelling approach are provided in Appendix B.
Figure 7: Graphical illustration of the food security risk index (median trajectories) for Egypt for the time period 2019-2050 under each SSP-RCP scenario and water-stress risk aggregation perspectives.
Combining the above projections we derive our estimations concerning food security under each scenario for Egypt and Ethiopia, in terms of the food security index introduced in Section 3.3 under four different perspectives (selection rules for the sensitivity parameter $\gamma$): (a) $\gamma=0$ (water stress is not taken into account), (b) very conservative perspective ($\gamma>0$ if $W_{c,t-1}>0.10$), (c) less conservative perspective ($\gamma>0$ if if $W_{c,t-1}>0.20$) and (d) no conservative perspective ($\gamma>0$ if if $W_{c,t-1}>0.40$). From the results in Figures 7, food security risk for Egypt seems to increase at the next decades, with higher rate at first years and with lower rate after 2040, under most of the SSP-RCP scenarios. Only for the SSP5-8.5 scenario after the year 2040 is indicated some small decrease on the risk. This picture is observed under all the different rules in selecting $\gamma$, except the case where the water risk component is omitted ($\gamma=0$). It is well known that Egypt is a highly water-stressed country and this picture is not expected to change in the near future. Without any doubt, availability of water resources is traditionally a critical issue for North Africa and especially Egypt, and it is still an unresolved problem. Therefore, the picture illustrated by the index for $\gamma=0$ is a quite unrealistic one, since the Egyptian agricultural sector strongly relies on water resources (see e.g. Christoforidou et al. (2023); Osman et al. (2016) and references therein) and therefore, food production activities are strongly related to water scarcity hazards mainly due to their current food production patterns, e.g. high emphasis in cereal production.
Figure 8: Graphical illustration of the food security risk index (median trajectories) for Ethiopia for the time period 2019-2050 under each SSP-RCP scenario and water-stress risk aggregation perspectives.
For the case of Ethiopia, the estimated food security risk under no aggregation with the water risk ($\gamma=0$), displays similar behaviour (Figure 8). Under all SSP-RCP scenarios, the food security risk seems to decline at a very high rate, almost exponentially. However, this case is much more different than Egypt. Although Ethiopia is not considered as a water-stressed country (the relevant water stress indicator is below 40%), over the following years this situation is expected to change dramatically due to the strictly increasing trend in the historical water stress records. According to our model for water stress, by the end of 2050 the median water stress for Ethiopia is projected to lie in the interval between 300% and 500% (please see Figure 13 in Appendix C) under all the scenarios considered! This means that Ethiopia should be able to acquire an extra amount of water resources corresponding to 3-5 times of the amount of water resources that are currently used. Such an occasion will significantly affect the food security risk for the country and unavoidably will affect the food production activities. Adopting the water stress penalization perspectives introduced in Section 3.3, a more realistic picture concerning the food security risk evolution in the next years is obtained. In particular, even under the NC perspective the food security risk is expected to greatly increase at next years under all SSP-RCP scenarios due to the forthcoming water scarcity hazard that is predicted by the water stress model. The same picture is observed under all perspectives (NC, LC and VC) indicating very significant food security risk to the end of 2050 and especially even in the middle scenario SSP2-4.5.
4.3 Food security risk estimates across SSP-RCPs
Although it is useful to obtain the food security risk estimate within each SSP scenario, this risk evaluation is characterised by the main drawback that it is not robust to the uncertainty concerning to which scenario materialises. Since the policy maker needs to take into account all possible scenarios in order to make a robust decision, we employ the relevant robust risk assessment approach discussed in Section 3.3. In terms of a toy example, we provide three different types of policy makers (or experts [^26] Please see the very interesting paper Sutherland and Burgman (2015) concerning the task of incorporating experts opinions): (a) one of complete ignorance, so all possible outcomes are weighted equally, (b) an optimistic one, where higher probability to scenarios more favourable for the environment is placed and (c) a pessimistic one, where higher probability to scenarios that are less favourable for the environment is allocated. These types are illustrated in Table 6.
Expert’s Opinion SSP1-1.9 SSP1-2.6 SSP2-4.5 SSP3-7.0 SSP4-6.0 SSP5-8.5
A. Ignorance 1/6 1/6 1/6 1/6 1/6 1/6
B. Optimistic 1/2 1/5 3/20 1/25 1/10 1/100
C. Pessimistic 1/100 1/25 1/10 1/5 3/20 1/2
Table 6: Realization probabilities of each SSP-RCP scenario according to different perspectives
Figure 9: Graphical illustration of the across scenario estimations of the food security index (left column) and water stress index (right column) for Egypt (upper panel) and Ethiopia (lower panel) for the time period 2019-2050.
In Figure 9 estimates for the food security index under the NC, LC and VC selection rules for $\gamma$ are illustrated across the SSP-RCP scenarios. For Egypt, the estimations under all perspectives concerning water stress almost coincide, and a quite homogeneous tendency for all expert opinions is displayed, indicating an increasing trend of concave type in food security risk until the year 2040, and then a small decrease. Clearly, this is a consequence of the high hazard concerning water scarcity that Egypt is exposed to. For the case of Ethiopia, the food security risk estimates are also very homogeneous, indicating increasing trend for food security risk under all perspectives and types of experts, evolving at almost exponentially rate as time progresses. Therefore, this robust approach in aggregating scenarios indicates that both countries are about to face important food security risks, however, the situation in Ethiopia is predicted to get worse at much higher pace unless proactive measures are taken. In any case, the provided across scenarios estimates for both countries, certify their robustness providing very homogeneous predictions which seem to be quite insensitive to alternations of the weights (different types of experts) concerning the considered set of SSP-RCP scenarios.
5 Conclusions
5.1 Limitations and Extensions
The various models employed in this paper are unavoidably subject to various assumptions and limitations. These are:
  • Raftery’s population model upon which our probabilistic scenario modelling was based, does not explicitly take into account population control policies or climate effects, but only implicitly through the past data on which the Bayesian model is trained, which may affect the outcomes of the model. The same comment is true for MaGE. In our modelling approach we introduce climate effects exogenously, through the use of RCP scenarios. The explicit introduction of such effects in the above models is clearly beyond the scope of the present work, and is an interesting subject for future extensions.
  • The data used for calibrating the model for food capacity were from the period before 2018 and therefore may not account for potential shocks of phenomena arising, such as e.g. pandemics or political events that happened after that. This is unavoidable due to the fact that such events are rather recent resulting to the unavailability of sufficient data for the period 2019-2023, during which these events happen. However, using techniques from extreme value theory the effects of these shocks may be assessed by stress testing the current model.
  • Clearly, the proposed methodology could be extended to construct a more detailed food security model incorporating detailed information on the infrastructure of the local food production structures, dietary patterns, etc. However, such a model would require collection of detailed data which for many countries are currently unavailable, at least in a reliable form. Therefore we chose to construct a rather aggregate model and a resulting food security risk index that relies on publicly available and easily accessible and reliable data.
5.2 Concluding Remarks
In this paper a general methodology for producing probabilistic socioeconomic scenarios compatible with the SSP-RCP framework and for assessing food security risk was proposed. The probabilistic scenarios represent the effects of the inherent uncertainty more efficiently than points estimates and are therefore better suited for projecting important socioeconomic quantities into the future. As an application of this methodology the issue of food security is considered, where a plausible index for its assessment is provided and a framework for the derivation of quantitative evaluations per scenario and future projections is proposed. While projections along scenarios are important, in practice there is uncertainty as to which scenario materializes. To address this question, a suitable framework for creating projections across scenarios using the concept of convex risk measures and their robust representation in the presence of model uncertainty is also proposed.
Our approach is illustrated by an application on assessing food security risk in two major countries in the upper Nile region, Egypt and Ethiopia. Examining each SSP-RCP scenario separately, our results show that under all possible trajectories, Egypt is not expected to face a serious food security risk for the next 30 years. On the contrary, although under all examined scenarios Ethiopia’s food security index shows an increasing trend, the pressure in country’s water resources is expected to rapidly and exponentially grow in the next years to levels that cannot be satisfied (about 3 to 5 times the country’s water resources would be needed till the year 2050). Such estimates should trigger the alarm for Ethiopia’s policy makers to design sustainable policies in terms of water use that should allow the country to avoid serious food security issues. The same conclusions carry over when we integrate all the above information, under three different hypothesis about the probability of scenario occurrence (i.e. complete ignorance, optimistic and pessimistic). Therefore, Ethiopian policy-makers should focus on possible pathways of food system transformation such as technology adoption, conflict reduction, strengthening resilience of the most vulnerable, lower the cost of nutritious foods, tackling poverty, etc. WHO et al. (2021) in an effort to mitigate the effects of the continuous food balance deficit, as well as of the sharpened food insecurity that is foreseen after 2030 due to rapidly increasing pressure in country’s water resources.
Acknowledgements
The authors would like to thank the Editor, the Associate Editor and the two anonymous referees for their insightful, constructive comments and suggestions that helped in improving the quality of the manuscript.
Funding
This project has received funding from the European Union’s Horizon 2020 partnership for research and innovation in the Mediterranean area programme (PRIMA) under grant agreement No 1942 (AWESOME).
Appendix A Extra Material from Section 2
Population evolution rates projections for Egypt and Ethiopia under the probabilistic setting
Figure 10: Relative population growth median projections and 90% zones for Egypt and Ethiopia under each SSP scenario using 2020 as the reference year.
RCP Scenarios Specifications
Scenario Emissions Level Temperature Change Mitigation Measures
RCP1.9 Best-case between $1-1.5^{o}$ Extremely stringent
RCP2.6 Low between $1.5-2^{o}$ Very stringent
RCP4.5 Medium - Low between $2.5-3^{o}$ Less stringent
RCP6.0 Medium - High between $3-3.5^{o}$ Very loose
RCP7.0 High up to $4^{o}$ until 2100 Extremely loose
RCP8.5 Worst-case up to $5^{o}$ until 2100 No mitigation
Table 7: Representative Concentration Pathways (RCP) definitions
Appendix B Uncertainty Quantification Issues
(a) Total population
(b) Minimum caloric requirements
(c) Food system capacity
(d) Level of water stress
(e) Food security risk (no water risk)
(f) Food security risk (NC)
(g) Food security risk (LC)
(h) Food security risk (VC)
Figure 11: Probabilistic projections for Egypt at the year 2050 for food security risk and relevant key quantities
(a) Total population
(b) Minimum caloric requirements
(c) Food system capacity
(d) Level of water stress
(e) Food security risk (no water risk)
(f) Food security risk (NC)
(g) Food security risk (LC)
(h) Food security risk (VC)
Figure 12: Probabilistic projections for Ethiopia at the year 2050 for food security risk and relevant key quantities
Appendix C Extra Material from Section 4
List of data and web sources used for the model fitting task
Train Dataset (Time Period: 1990 - 2018)
Driver Unit Data Source
Agricultural land 1000 ha FAO Database
GDP (per capita) USD (constant 2015) World Bank Database
Labour (total) 1000 people World Bank Database
Labour (in agricultural activities) % of total labour World Bank Database (ILO estimates)
Population 1000 people World Bank Database
Food Supply Kcal/capita/day FAO Database
Food Production Quantity tonnes FAO Database
Exported food quantity tonnes FAO Database
Imported food quantity tonnes FAO Database
Precipitation (yearly average) mm Climate Knowledge Portal
Temperature (yearly average) degrees of Celsium Climate Knowledge Portal
Water Stress Indicator (SDG 6.4.2) % of internal water resources FAO/AQUASTAT Database
Table 8: Online data sources from which data where used for tuning the models for food system capacity stated in Section 3.2
Cobb-Douglas Modelling Approach
For each country the following set of parametric models (Cobb-Douglas type for both layers) are considered and fitted to data:
(a) Set of Upper Layer Models
$\left\{\begin{array}[]{ll}C_{c,t}^{FS}\sim&\alpha_{C,0}e^{\alpha_{C,1}t}\left(Q_{c,t}^{Dom}\right)^{\alpha_{C,2}}\left(Q^{Exp}_{c,t}\right)^{\alpha_{C,3}}\left(Q^{Imp}_{c,t}\right)^{\alpha_{C,4}}\\ Q_{c,t}^{Dom}\sim&\alpha_{D,0}e^{\alpha_{D,1}t}\,\,\left(P_{c,t}\right)^{\alpha_{D,2}}\left(I_{c,t-1}\right)^{\alpha_{D,3}}\left(L_{c,t}^{Agr}\right)^{\alpha_{D,4}}\left(W_{c,t}\right)^{\alpha_{D,5}}\left(A_{c,t}\right)^{\alpha_{D,6}}\left(T_{c,t}\right)^{\alpha_{D,7}}\\ W_{c,t}\sim&\alpha_{W,0}e^{\alpha_{W,1}t}\left(P_{c,t}\right)^{\alpha_{W,2}}\left(I_{c,t-1}\right)^{\alpha_{W,3}}\left(Q^{Dom}_{c,t}\right)^{\alpha_{W,4}}\left(A_{c,t}\right)^{\alpha_{W,5}}\left(T_{c,t}\right)^{\alpha_{W,6}}\left(Pr_{c,t}\right)^{\alpha_{W,7}}\\ A_{c,t}\sim&\alpha_{A,0}e^{\alpha_{A,1}t}\,\,\left(P_{c,t}\right)^{\alpha_{A,2}}\left(I_{c,t-1}\right)^{\alpha_{A,3}}\left(Q_{c,t-1}^{Dom}\right)^{\alpha_{A,4}}\\ \end{array}\right.$
(b) Set of Lower Layer Models
$\left\{\begin{array}[]{ll}Q^{Exp}_{c,t}\sim&\beta_{E,0}e^{\beta_{E,1}t}\,\,\left(P_{c,t}\right)^{\beta_{E,2}}\left(I_{c,t-1}\right)^{\beta_{E,3}}\\ Q^{Imp}_{c,t}\sim&\beta_{I,0}e^{\beta_{I,1}t}\,\,\left(P_{c,t}\right)^{\beta_{I,2}}\left(I_{c,t-1}\right)^{\beta_{I,3}}\\ L^{Agr}_{c,t}\sim&\beta_{L,0}e^{\beta_{L,1}t}\,\,\left(P_{c,t}\right)^{\beta_{L,2}}\left(I_{c,t-1}\right)^{\beta_{L,3}}\left(L_{c,t}\right)^{\beta_{L,4}}\end{array}\right.$
following the abbreviations defined in Section 3.2, for $c=EGY,ETH$ with country-specific parameter vectors ${\bm{\alpha}}_{c}:=({\bm{\alpha}}_{C,c},{\bm{\alpha}}_{D,c},{\bm{\alpha}}_{W,c},{\bm{\alpha}}_{A,c})^{\prime}$ and ${\bm{\beta}}_{c}:=({\bm{\beta}}_{E,c},{\bm{\beta}}_{I,c},{\bm{\beta}}_{L,c})^{\prime}$ for the upper and the lower layer models, respectively.
The estimated models for Egypt and Ethiopia
Upper Layer Lower Layer
Predictor
Food System Capacity
Water Stress
Dom. Food Prod. Qty
Crop Land (used)
Labour (Agr)
Exported Food Qty
Imported Food Qty
Technology coef.
68.4069
  0.2254
  5.6151
0.9429
2.8004
0.0011
0.4020
Yearly trend
  0.01600.0160\,\,0.01600.0160
-0.0010
  0.0045
0.0022
0.0030
0.0296
0.0124
Population
-0.0799
  0.2275
0.1131
0.0580
1.3062
0.5991
Labour (Total)
0.1580
Labour (Agr)
  0.1957
GDP (per capita)
-0.0705
  0.1202
0.0299
0.0676
0.8897
0.5627
Exported Qty
-0.0149
Imported Qty
  0.03100.0310\,\,0.03100.0310
Dom. Food Prod.
  0.24190.2419\,\,0.24190.2419
  0.2065
0.1873
Water Stress
  0.3373
Crop Land (used)
  0.3869
  0.4858
Temperature (avg)
  0.3480
-0.0483
Precipitation (avg)
-0.0695
Explained Deviance ($R^{2}$)
99.53%
  44.56%
94.27%
89.21%
58.11%
92.70%
89.48%
Table 9: The two-layer model parameter estimates for Egypt using as trainset the available data from period 1990-2018
Upper Layer Lower Layer
Predictor
Food System Capacity
Water Stress
Dom. Food Prod. Qty
Crop Land (used)
Labour (Agr)
Exported Food Qty
Imported Food Qty
Technology coef.
58.4825
  0.0039
  0.0005
  0.1796
  1.1033
0.0127
0.0889
Yearly trend
0.0272
  0.0128
  0.0177
  0.0224
  0.0054
0.0235
0.0199
Population
  0.5546
  0.6245
  1.1185
  0.2827
0.8926
0.6915
Labour (Total)
  0.5119
Labour (Agr)
  0.4972
GDP (per capita)
-0.2472
  0.2818
-0.0599
-0.1051
0.5334
0.5877
Exported Qty
0.0181
Imported Qty
0.0198
Dom. Food Prod.
0.2795
-0.0900
-0.3790
Water Stress
-0.1111
Crop Land (used)
  1.4905
  0.2077
Temperature (avg)
  1.1754
  1.7254
Precipitation (avg)
-0.8630
Explained Deviance ($R^{2}$)
99.78%
95.74%
97.80%
90.21%
99.62%
73.62%
64.14%
Table 10: The two-layer model parameter estimates for Ethiopia using as trainset the available data from period 1990-2018
List of data and models used for the projection task
Projections Data Sources & Models
Driver Unit Scenario Dependence Data Source/model
GDP (per capita) USD (constant 2015) SSP MaGE
Labour (total) 1000 people SSP MaGE
Population 1000 people SSP BayesPop + SSP Modeller
Precipitation (yearly average) mm RCP Climate Knowledge Portal
Temperature (yearly average) degrees of Celsium RCP Climate Knowledge Portal
Table 11: Data sources and models used for creating projections on basic socioeconomic and climate drivers for the time period 2019-2050
Water-stress projections under SSP-RCP scenarios
Figure 13: Graphical illustration of the water stress index for Egypt and Ethiopia for the time period 2019-2050 under each scenario (median estimates).